# On the Complexity of Determinations

### 30 Mar 2026 v1

@@ 3.3 Distributed Extension and Pointer Chasing
Setting $b_{\ell}=0$ recovers the width bound; setting $w=1$ gives a communication lower bound. Both are tight. Proof in Appendix C.2.
@@
At sparsity $s=1$ the specification is functional (determination depth 0); hardness appears only when constraints are distributed ($k$-step pointer chasing, $\Omega(n/k)$-bit tradeoff [NW93]). At $s\geq 2$ the specification is relational (determination depth $k$); the exponential width bound applies even to a centralized machine. The transition from polynomial to exponential occurs exactly at $s=1$ vs. $s\geq 2$—the boundary between functions and relations.
@@ 4 Structural Results
The exponential separation shows that determination depth is a meaningful complexity measure with concrete consequences. We now establish structural results that explain why the separation works and what it implies for the general theory. The results in this section concern the interaction between determination depth and computational depth; communication re-enters in the distributed setting (Appendix B). The conservation law (Theorem 3) shows that enriching the commitment basis trades determination layers for computational depth within layers, but the total sequential cost is bounded below—explaining the tightness witnessed by the constrained generation task. The oracle characterization (Proposition 1) shows that under a commutative basis in the online setting, even a strategy with access to an unbounded disclosure oracle cannot reduce the number of layers below $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$. Finally, stable matching (Proposition 2) witnesses the orthogonality of the two axes: computational cost is polynomial regardless of determination depth.
@@ 4.1 Conservation of sequential depth
A specification may have internal dependency structure—a chain of commitments where each depends on the previous one’s outcome—even in the offline setting, where forward validity constraints are absent. The following theorem says this structure cannot be eliminated.
@@
A richer basis can reduce the number of layers $d$ by bundling multiple constant-depth commitments into a single commitment, but evaluating that commitment requires circuit depth $c_{i}$ that accounts for the bundled dependencies. The bound says the total sequential depth of commitment evaluation—summing $(1+c_{i})$ over layers—cannot fall below $\mathsf{Depth}_{\Phi_{0}}(\mathsf{Spec})$. The bound is not a counting identity: a richer basis could in principle exploit algebraic cancellations to resolve the specification in fewer total sequential steps than the longest dependency chain, and the theorem’s content is that no such shortcut exists. The constrained generation task (Section 3) witnesses tightness: for every split of the $k$ dependency links, the bound is achieved with equality (Appendix C.2). The proof (Appendix C.1) traces the longest path in the $\Phi_{0}$-dependency DAG: each pair either crosses a layer boundary (contributing to $d$) or forces a circuit path within a layer (contributing to $c_{i}$).
@@ 4.2 Oracle characterization
The conservation law bounds the sequential depth of commitment evaluation in both the offline and online settings. What if the strategy itself has unbounded computational power between commitments? In the online setting, this question is non-trivial:
@@
Under a commutative basis $\Phi$ in the online setting, grant the strategy a free disclosure oracle call at the start of every layer (a function that may inspect the entire history, perform unbounded computation, and return any value). Any determination strategy that resolves $\mathsf{Spec}$ still requires at least $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ layers.
@@
The conservation law and the oracle characterization are complementary: the former bounds computation within layers (commitment evaluation), the latter bounds the number of layers themselves (strategy computation between layers cannot reduce it). In the distributed setting, the oracle characterization serves as a single-agent baseline: Appendix B shows that asynchronous multi-agent systems cannot always achieve this baseline, recovering the Halpern–Moses impossibility of common knowledge as a special case and providing quantitative depth predictions beyond the binary threshold.
4.3 Orthogonality: stable matching
Within a single specification, the conservation law couples determination and computational depth. Across specifications, these are independent: neither subsumes the other. Stable matching witnesses this sharply.
Given preference lists for two disjoint sets of $n$ agents, the specification maps the input to the set of all stable matchings [GS62]. The structural theory of stable matchings is organized around rotations—cyclic reassignments that transform one stable matching into an adjacent one in the lattice of stable matchings [IL86, GI89]. Rotations are partially ordered by precedence ($\rho$ must be applied before $\rho^{\prime}$ can be exposed); the resulting rotation poset encodes the determination structure.
Proposition 2 (Stable matching depth).
For any stable matching instance, determination depth under the rotation basis equals the height of the rotation poset [IL86, GI89]. By the Irving–Leather realization theorem, every finite poset arises as a rotation poset, so every determination depth is achievable.
Finding a stable matching is in P [GS62], yet determination depth can be arbitrarily large: computational cost is polynomial regardless of determination depth. The Gale–Shapley algorithm finds a matching in $O(n^{2})$ steps but is not depth-optimal; a strategy that applies all exposed rotations simultaneously at each layer achieves depth equal to the poset height, recovering the parallel algorithm of Garg [GAR20]. Proof and full development in Appendix D.1.
@@ 5 Metacomplexity of Determination Depth
The polynomial hierarchy is thus precisely the hierarchy of determination depths for circuit-encoded specifications. (Offline, the metacomplexity is NP-hard but does not capture the full hierarchy; see Proposition 3.) Where circuit depth and communication complexity parameterize the PH through computational resources, determination depth parameterizes it through a semantic one—layers of irrevocable choice. The correspondence is not the contribution; it is a calibration. An alternating Turing machine has a fixed quantifier prefix determined by the program; here the determiner optimizes which variables to control and in which order, so the quantifier schedule is itself part of the problem. More fundamentally, the paper’s main results—the exponential separation (Theorem 1) and the conservation law (Theorem 3)—are offline results about depth–width and depth–computation tradeoffs within a single specification, a regime not captured by the quantifier alternation that ATMs model.
@@ 6 Conclusion
Complexity theory has measured the cost of computing a uniquely determined answer. This paper introduces determination depth to measure the cost of committing to an answer when many are admissible—a semantic complexity measure orthogonal to computation. The exponential separation (Theorem 1) shows this cost is real, the three-way tradeoff (Theorem 2) shows depth, width, and communication are fungible, the PH characterization (Theorem 4) shows it is well-calibrated against the classical hierarchy. In the distributed setting, the framework recovers the Halpern–Moses impossibility of asynchronous common knowledge as a special case (Theorem 5), and the conservation law provides quantitative depth–cost predictions beyond the binary threshold. The appendices ground the framework in BSP round complexity, chain-of-thought reasoning, stable matching, extensive-form games, and distributed graph coloring; Appendix F surveys related work in detail. A number of directions remain open:
@@
Provenance of determinations. Classical data provenance explains query answers via a commutative semiring over monotone derivations [GKT07]. Determinations introduce a different algebra: a conditionally commutative monoid of layered commitments. The two structures do not align—semiring provenance applies only after a determination has resolved all ambiguity, and cannot represent how alternative commitments would have led to different explanations. A unifying algebraic theory of determination provenance— tracking how commitments at each layer shape downstream explanations—remains open.
@@ A.1 Commuting Layers and Depth
We extend the $\cdot$ notation to multisets: $H\cdot L$ is the history obtained by appending all commitments in $L$ after $H$ (in any order).
@@
A finite multiset $L$ of commitments from a common basis $\Phi$ is a commuting layer for $\mathsf{Spec}$ at $H$ if applying its commitments after $H$ is order-independent: for any two listings $\psi_{1},\ldots,\psi_{t}$ and $\psi^{\prime}_{1},\ldots,\psi^{\prime}_{t}$ of the elements of $L$ (respecting multiplicity),
@@
Let $D(H)=\varphi_{1}\cdot\cdots\cdot\varphi_{m}$ be the determination of a history $H$ over $\Phi$. A run is a maximal subsequence of consecutive commitment events in $H$ (with no intervening environment events). A layering of a run $R=\psi_{1}\cdot\cdots\cdot\psi_{r}$ is a partition into nonempty multisets $L_{1},\ldots,L_{k}$ such that each $L_{i}$ is a commuting layer for $\mathsf{Spec}$ at the history $H_{0}\cdot L_{1}\cdots L_{i-1}$, where $H_{0}$ is the history prefix immediately before the run. The depth of a run is the minimum $k$ over all its layerings. The depth of $D(H)$, written $0pt(D(H))$, is the sum of the depths of its runs. In the offline setting, the entire determination is a single run.
For a single run, we write $L_{1}\mathbin{\triangleright}L_{2}\mathbin{\triangleright}\cdots\mathbin{\triangleright}L_{k}$ to denote the layered determination; the $\mathbin{\triangleright}$ notation makes the layering (and hence depth) visible. Every run admits a depth-minimal layering; in the offline setting (a single run), we call this the layered normal form. Depth is bounded by cost ($0pt(D)\leq\mathsf{cost}(D)$, since layers are nonempty), but can be much smaller when most commitments commute.
@@ A.2 Dependency Chains and Depth Characterization
The offline results (Proposition 4 below) use local forced dependency. The distributed results (Theorem 5) use universal forced dependency. In the online setting, depth may exceed the longest dependency chain: run boundaries force additional layers even between independent commitments. The oracle characterization (Proposition 1) handles the online setting directly.
@@ Appendix B Distributed Determinations and Common Knowledge
We extend the framework to multiple agents. Fix a finite set of agents $\{1,\ldots,n\}$. Each event in a history $H$ is associated with an agent; for agent $p$, the projection $H|_{p}$ is the subgraph of $H$ induced on $p$’s events—local computation steps, message sends, and message receives at pp—with only the edges of $H$ whose both endpoints are events at $p$. (Causal paths through other agents’ events are not visible in the projection; this is the standard Lamport projection.) The partial order on $H|_{p}$ is the transitive closure of $\to$ restricted to $p$’s events: $e_{1}$ precedes $e_{2}$ in $H|_{p}$ whenever $e_{1}\to^{*}e_{2}$ in $H$ and both are events at $p$. The frontier of $H|_{p}$ is the set of sinks of $H|_{p}$ (events with no successors under $\to$).
@@
At a synchronization point, every agent sees the same admissible set (condition (iii)), and each agent knows the synchronization occurred (condition (i): $e^{*}_{j,p}$ is in its projection).
@@
Assume the synchronization protocol (conditions (i)–(iii)) is common knowledge among all agents. Then at each synchronization point, the current admissible set is common knowledge.
@@
Condition (iii) gives $E(S,H)$; common knowledge of the protocol iterates this to all levels. ∎
@@
The oracle characterization (Section 4.2) assumes a single global oracle that sees the entire history. In the distributed setting, each agent has access only to a local oracle that sees its projection $H|_{p}$. A synchronization round (conditions (i)–(iii)) is precisely the mechanism that elevates local oracles to global ones: after synchronization, every agent’s local state includes the same admissible set (by condition (iii) and Lemma 1), so a local oracle can determine the admissible set without global access. Between synchronization points, local oracles are strictly weaker—each sees only its own projection.
@@
Appendix C Proofs for Conservation and Tightness
C.1 Conservation of sequential depth (Theorem 3)
Let $\Phi_{0}$ be a constant-depth basis for $\mathsf{Spec}$ (Definition 11), and let $G_{\Phi_{0}}$ be its forced-dependency DAG (Definition 20). Each edge $\varphi\to\psi$ in $G_{\Phi_{0}}$ has a concrete witness exclusion: a triple $(H,H\cdot\varphi,o)$ such that excluding outcome $o$ at $H$ is invalid, but excluding $o$ at $H\cdot\varphi$ is valid—the dependency is witnessed by a specific outcome that cannot be excluded until $\varphi$ has taken effect. Any basis $\Phi$ that resolves $\mathsf{Spec}$ must also exclude each witness outcome $o$; the conservation law measures how $\Phi$ pays for these exclusions.
Observation 4 (Semantic dependency implies circuit dependency).
Let $\varphi_{j}$ and $\varphi_{j+1}$ be commitments from a constant-depth basis $\Phi_{0}$ such that $\varphi_{j+1}$ is forced to depend on $\varphi_{j}$. If a circuit $C$ computes a valid layer of a determination that achieves a witness exclusion of both $\varphi_{j}$ and $\varphi_{j+1}$, then $C$ contains a directed path from the gates computing $\varphi_{j}$’s witness exclusion to the gates computing $\varphi_{j+1}$’s witness exclusion.
Proof.
By forced dependency, there exist instances where $\varphi_{j+1}$’s exclusion is invalid unless $\varphi_{j}$’s exclusion has already occurred. If $C$ contained no path from $\varphi_{j}$’s gates to $\varphi_{j+1}$’s, then $\varphi_{j+1}$’s output would be independent of $\varphi_{j}$’s, and on such an instance the circuit would produce an invalid exclusion—a contradiction. ∎
Theorem 3 (Conservation law, restated). For any basis $\Phi$ and any valid determination over $\Phi$ that resolves $\mathsf{Spec}$ in $d$ layers with per-layer circuit depths $c_{1},\ldots,c_{d}$, $\sum_{i=1}^{d}(1+c_{i})\geq d^{*}$, where $d^{*}=\mathsf{Depth}_{\Phi_{0}}(\mathsf{Spec})$ is the determination depth under a constant-depth basis $\Phi_{0}$.
Proof of Theorem 3.
Let $\varphi_{1}\to\cdots\to\varphi_{p}$ be a longest path in $G_{\Phi_{0}}$, so $p=d^{*}$ (Corollary 4). Each edge $\varphi_{j}\to\varphi_{j+1}$ has a witness exclusion $(H_{j},H_{j}\cdot\varphi_{j},o_{j})$: outcome $o_{j}$ cannot be excluded at $H_{j}$ until $\varphi_{j}$ has taken effect. Fix any determination $D$ under basis $\Phi$ with $d$ layers and per-layer circuit depths $c_{1},\ldots,c_{d}$. Since $D$ resolves $\mathsf{Spec}$, each $o_{j}$ is eventually excluded; let $\lambda(j)$ be the $\Phi$-layer in which $o_{j}$ is first excluded. The richer basis $\Phi$ may exclude multiple witness outcomes within a single layer.
For each consecutive pair $\varphi_{j}\to\varphi_{j+1}$ in the path ($j=1,\ldots,p\!-\!1$), exactly one of the following holds.
Case 1: $\lambda(j)<\lambda(j+1)$. The predecessor’s effect is achieved in a strictly earlier $\Phi$-layer. This pair contributes at least $1$ to $d$.
Case 2: $\lambda(j)=\lambda(j+1)=i$. Both effects are achieved in the same $\Phi$-layer $i$. By Observation 4, the circuit computing $\Phi$-layer $i$’s commitments contains a directed path from the gates for $\varphi_{j}$’s exclusion to those for $\varphi_{j+1}$’s, contributing at least $1$ to $c_{i}$.
Case 3: $\lambda(j)>\lambda(j+1)$. The successor’s effect is achieved in an earlier $\Phi$-layer than the predecessor’s. Then $\varphi_{j+1}$’s exclusion is valid without $\varphi_{j}$’s prior exclusion, contradicting forced dependency. Hence Case 3 does not arise.
Each of the p1p-1 consecutive pairs contributes at least $1$ to either $d$ (Case 1: a layer boundary) or some $c_{i}$ (Case 2: a circuit path within a layer), and each layer visited by the path contributes its $1$ term. Hence $\sum_{i=1}^{d}(1+c_{i})\geq p=d^{*}$. ∎

### 01 Apr 2026 v2

@@ 3.3 Distributed Extension and Pointer Chasing
Setting $b_{\ell}=0$ recovers the width bound; setting $w=1$ gives a communication lower bound. Both are tight. Proof in Appendix C.
@@
At sparsity $s=1$ the specification is functional (determination depth 0); the classical round complexity of $k$-step pointer chasing [NW93] arises entirely from distributing the chain across agents, not from relational choice. At $s\geq 2$ the specification is relational (determination depth $k$); the exponential width bound applies even to a centralized machine. The transition occurs exactly at $s=1$ vs. $s\geq 2$—the boundary between functions and relations.
@@ 4 Structural Results
The exponential separation shows that determination depth is a meaningful complexity measure with concrete consequences. We now establish structural results that explain why the separation works and what it implies for the general theory. The sequential cost of determination is governed by the dependency structure among commitments—a structure that computation cannot alter. The conservation law (Theorem 3) shows that enriching the commitment basis merely relabels determination layers as circuit depth within layers—the total sequential cost is conserved. The oracle characterization (Proposition 1) shows that under a commutative basis in the online setting, even unbounded computation between layers cannot reduce their number. Communication re-enters in the distributed setting (Appendix B).
@@ 4.1 Conservation of sequential depth
A specification may have internal dependency structure—a chain of commitments where each depends on the previous one’s outcome—even in the offline setting, where forward validity constraints are absent. A richer basis can reduce the number of determination layers $d$ by bundling multiple constant-depth commitments into a single commitment, but evaluating that commitment requires circuit depth $c_{i}$ that accounts for the bundled dependencies. The total sequential depth of commitment evaluation—summing $(1+c_{i})$ over layers—cannot fall below $\mathsf{Depth}_{\Phi_{0}}(\mathsf{Spec})$. This is not a counting identity: a richer basis could in principle exploit algebraic cancellations to resolve the specification in fewer total sequential steps than the longest dependency chain, and the following theorem’s content is that no such shortcut exists.
@@
Each edge $\varphi_{j}\to\varphi_{j+1}$ in the forced-dependency DAG of $\Phi_{0}$ has a witness exclusion: a triple $(H_{j},H_{j}\cdot\varphi_{j},o_{j})$ such that excluding $o_{j}$ at $H_{j}$ is invalid, but excluding $o_{j}$ at $H_{j}\cdot\varphi_{j}$ is valid. The proof traces the longest path in this DAG and shows that each edge must be “paid for” by either a layer boundary or a circuit path within a layer. The key step is:
Observation 4 (Semantic dependency implies circuit dependency).
If $\varphi_{j+1}$ is forced to depend on $\varphi_{j}$ (under a constant-depth basis $\Phi_{0}$), and a circuit $C$ computes a layer that achieves the witness exclusions of both, then $C$ contains a directed path from the gates for $\varphi_{j}$’s exclusion to those for $\varphi_{j+1}$’s.
Proof.
By forced dependency, $\varphi_{j+1}$’s exclusion is invalid unless $\varphi_{j}$’s has occurred. If $C$ had no such path, $\varphi_{j+1}$’s output would be independent of $\varphi_{j}$’s. But the witness exclusion guarantees instances where excluding $o_{j+1}$ without first excluding $o_{j}$ is invalid, so $C$ would produce an invalid refinement on such an instance—a contradiction. ∎
Proof of Theorem 3.
Let $d^{*}=\mathsf{Depth}_{\Phi_{0}}(\mathsf{Spec})$ and let $\varphi_{1}\to\cdots\to\varphi_{d^{*}}$ be a longest forced-dependency path in the $\Phi_{0}$-dependency DAG (Proposition 3). Each edge $\varphi_{j}\to\varphi_{j+1}$ has a witness exclusion. Fix any determination under basis $\Phi$ with $d$ layers and per-layer circuit depths $c_{1},\ldots,c_{d}$. Since the determination resolves $\mathsf{Spec}$, each $o_{j}$ is eventually excluded; let $\lambda(j)$ be the $\Phi$-layer in which $o_{j}$ is first excluded from the admissible set at all extensions.
For each consecutive pair $\varphi_{j}\to\varphi_{j+1}$, exactly one of the following holds. Case 1: $\lambda(j)<\lambda(j+1)$—the pair crosses a layer boundary, contributing at least $1$ to $d$. Case 2: $\lambda(j)=\lambda(j+1)=i$—both exclusions occur in the same layer; by forced dependency, the circuit computing layer $i$ must contain a directed path from the gates for $o_{j}$’s exclusion to those for $o_{j+1}$’s, contributing at least $1$ to $c_{i}$. Case 3: $\lambda(j)>\lambda(j+1)$—the successor is excluded before the predecessor, contradicting forced dependency.
Each of the $d^{*}-1$ pairs contributes at least $1$ to either $d$ or some $c_{i}$, and each layer visited contributes its $1$ term. Hence $\sum_{i=1}^{d}(1+c_{i})\geq d^{*}$. ∎
The constrained generation task (Section 3) witnesses tightness: for every split of the $k$ dependency links, the bound is achieved with equality (Appendix C).
@@ 4.2 Oracle characterization
The conservation law shows that enriching commitments cannot reduce the total sequential cost. What if instead the strategy has unbounded computational power between commitments? In the online setting, this question is non-trivial:
@@
Under a commutative basis $\Phi$ in the online setting, grant the strategy a free call to a disclosure oracle at the start of every layer (a function that may inspect the entire history, perform unbounded computation, and return any value). Any determination strategy that resolves $\mathsf{Spec}$ still requires at least $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ layers.
@@
4.3 Relationship to computational complexity
Determination depth measures a cost that classical complexity does not track. A function has determination depth 0 regardless of its circuit depth; a relation can require exponential width to compensate for any reduction in depth, even when every individual commitment is a constant-time operation (Theorem 1). This width cost is purely relational: at sparsity $s=1$ (a function), it vanishes entirely (Observation 3). In the online setting, the oracle characterization shows that determination depth is irreducible: no amount of computation between layers can reduce their number.
Classical computational complexity plays a narrow role within the framework: each commitment is itself a function, so its evaluation has classical circuit depth. The conservation law shows that this is all computation can do—it emulates atomic-basis determination layers one-for-one as circuit layers within a richer commitment, preserving the total sequential cost exactly. Determination depth under the atomic basis is thus a universal lower bound on the total sequential cost of resolution, regardless of the basis used.
Determination depth on the atomic basis therefore provides a complete accounting of the sequential cost of resolving a relation—both commitment dependency and forward validity, in both offline and online settings. Computational depth captures only the cost of evaluating individual commitments, a sub-problem that arises when the basis is enriched and only in settings where forward validity is absent.
In the distributed setting, the oracle characterization serves as a single-agent baseline: Appendix B shows that asynchronous multi-agent systems cannot always achieve this baseline, recovering the Halpern–Moses impossibility of common knowledge as a special case and providing quantitative depth predictions beyond the binary threshold.
@@ 5 Metacomplexity of Determination Depth
The polynomial hierarchy is thus precisely the hierarchy of determination depths for circuit-encoded specifications. (Offline, the metacomplexity is NP-hard but does not capture the full hierarchy; see Proposition 2.) Where circuit depth and communication complexity parameterize the PH through computational resources, determination depth parameterizes it through a semantic one—layers of irrevocable choice. The correspondence is not the contribution; it is a calibration. An alternating Turing machine has a fixed quantifier prefix determined by the program; here the determiner optimizes which variables to control and in which order, so the quantifier schedule is itself part of the problem. The alternation arises from the interaction between commitments and environment extensions, not from the encoding of the specification. More fundamentally, the paper’s main contributions—the exponential separation (Theorem 1), the conservation law (Theorem 3), and the oracle characterization (Proposition 1)—establish that determination depth resists computation in both the offline and online settings, a phenomenon not captured by quantifier alternation.
@@ 6 Conclusion
Complexity theory has measured the cost of computing a uniquely determined answer. This paper introduces determination depth to measure the cost of committing to an answer when many are admissible—a cost that no amount of computation can eliminate. The exponential separation (Theorem 1) shows this cost is real, the conservation law (Theorem 3) and oracle characterization (Proposition 1) show that computation cannot reduce it, and the PH characterization (Theorem 4) shows it is well-calibrated against the classical hierarchy. In the distributed setting, the framework recovers the Halpern–Moses impossibility of asynchronous common knowledge as a special case (Appendix B), and the depth–width–communication tradeoff (Theorem 2) shows that the depth, width, and communication costs of resolving a specification are fungible. The appendices ground the framework in BSP round complexity, chain-of-thought reasoning, stable matching, extensive-form games, and distributed graph coloring; Appendix F surveys related work in detail. A number of directions are open:
@@
Provenance of determinations. Classical data provenance explains query answers via a commutative semiring over monotone derivations [GKT07]. Determinations introduce a different algebra: a conditionally commutative monoid of layered commitments. The two structures do not align naturally. A unifying algebraic theory of determination provenance— tracking how commitments at each layer shape downstream explanations—remains open.
@@ A.1 Commuting Layers and Depth
We extend the $\cdot$ notation to sets: $H\cdot L$ is the history obtained by appending all commitments in $L$ after $H$ (in any order).
@@
A finite set $L$ of commitments from a common basis $\Phi$ is a commuting layer for $\mathsf{Spec}$ at $H$ if applying its commitments consecutively after $H$ (with no intervening environment events) is order-independent: for any two listings $\psi_{1},\ldots,\psi_{t}$ and $\psi^{\prime}_{1},\ldots,\psi^{\prime}_{t}$ of the elements of $L$,
@@
Let $D(H)=\varphi_{1}\cdot\cdots\cdot\varphi_{m}$ be the determination of a history $H$ over $\Phi$. A run is a maximal subsequence of consecutive commitment events in $H$ (with no intervening environment events). A layering of a run $R=\psi_{1}\cdot\cdots\cdot\psi_{r}$ is a partition into nonempty sets $L_{1},\ldots,L_{k}$ such that each $L_{i}$ is a commuting layer for $\mathsf{Spec}$ at the history $H_{0}\cdot L_{1}\cdots L_{i-1}$, where $H_{0}$ is the history prefix immediately before the run. The depth of a run is the minimum $k$ over all its layerings. The depth of $D(H)$, written $0pt(D(H))$, is the sum of the depths of its runs. In the offline setting, the entire determination is a single run.
Every run admits a depth-minimal layering; in the offline setting (a single run), we call this the layered normal form. Depth is bounded by cost ($0pt(D)\leq\mathsf{cost}(D)$, since layers are nonempty), but can be much smaller when most commitments commute.
@@ A.2 Dependency Chains and Depth Characterization
The offline results (Proposition 3 below) use local forced dependency. The distributed results (Theorem 5) use universal forced dependency. In the online setting, the oracle characterization (Proposition 1) handles depth directly, without reducing to dependency chains.
@@ Appendix B Distributed Determinations and Common Knowledge
We extend the framework to multiple agents. Fix a finite set of agents $\{1,\ldots,n\}$. Each event in a history $H$ is associated with an agent; for agent $p$, the projection $H|_{p}$ is the restriction of $H$ to $p$’s events—local computation steps, message sends, and message receives at pp—ordered by the transitive closure of $\to$ restricted to $p$’s events: $e_{1}$ precedes $e_{2}$ in $H|_{p}$ whenever $e_{1}\to^{*}e_{2}$ in $H$ and both are events at $p$. (Causal paths through other agents’ events are not directly visible, but their ordering effects are preserved.) The frontier of $H|_{p}$ is the set of sinks of $H|_{p}$ (events with no successors under $\to$).
@@
At a synchronization point, every agent sees the same admissible set (condition (ii)), and each agent knows the synchronization occurred ($e^{*}_{j,p}$ is in its projection).
@@
Assume the synchronization-point conditions (i)–(ii) are common knowledge among all agents. Then at each synchronization point, the current admissible set is common knowledge.
@@
Condition (ii) gives $E(S,H)$; common knowledge of the conditions iterates this to all levels. ∎
@@
The oracle characterization (Section 4.2) assumes a single global oracle that sees the entire history. In the distributed setting, each agent has access only to a local oracle that sees its projection $H|_{p}$. A synchronization round (conditions (i)–(ii)) is precisely the mechanism that elevates local oracles to global ones: after synchronization, every agent’s local state includes the same admissible set (by condition (ii) and Lemma 1), so a local oracle can determine the admissible set without global access. Between synchronization points, local oracles are strictly weaker—each sees only its own projection.
@@
Appendix C Tightness Proofs